Lance Whittingham IV specializes in buying deep discount bonds. These represent bonds that are trading at well below par value. He has his eye on a bond issued by the Leisure Time Corporation. The $1,000 par value bond pays 6 percent annual interest and has 15 years remaining to maturity. The current yield to maturity on similar bonds is 15 percent. UseAppendix B and Appendix D for an approximate answer but calculate your final answer using the formula and financial calculator methods.

a. What is the current price of the bonds? (Do not round intermediate calculations. Round your final answer to 2 decimal places. Assume interest payments are annual.)

b. By what percent will the price of the bonds increase between now and maturity? (Do not round intermediate calculations. Input your answer as a percent rounded to 2 decimal places.)
Answer #1

Solution to A

Given,

Face value of bond = $1,000 per bond

yield-to-maturity = 15%

The bond expires in 15 years

Coupon payment is 6% annually or 0.06 *$1000 or $60

Input these variables in a financial calculator to compute current price of the bond. The current value of the bond is $473.73

If you don't have a financial calculator, you can use the following formula to calculate the bond price.

((coupon payment at year 1)/(1 + yield)^1) + (coupon payment at year 2)/(1 + yield)^2 .... + (coupon payment at year 15 + face value of bond)/(1+yield)^15

= (60/1.15) + (60/ (1.15)^2) .... + (1060/ 1.15^15)

Basically, you need to bring back the future coupon payments and the final principal payment, back in present terms.

= $52.17 + $45.37 ..... +$130.27

=$473.73

Solution to B

The price of the bond reaches face value when it comes close to maturity. At maturity, the price of the bond is the face value, as that is the only payment remaining, and we don't have to discount that.

Hence the appreciation between now and maturity will be ((face value of the bond)/ (current value of the bond) -1) *100

= (1,000/473.73-1)*100 = 111.09%